Given the wave equation y(x, t) = 2.0 Sin(3.0x – 4.0t + π

2

)m, where x is


in meter and t is in second. Find the:

a. Amplitude

b. Frequency

c. Wavelength

d. Speed

e. Initial height at x=1.0m


1
Expert's answer
2022-01-06T09:06:44-0500

Given:


"y(x, t) = 2.0 \sin(3.0x – 4.0t + π/2)"

The general form of a wave is given by

"y(x, t) = A \sin
\left(kx – \omega t +\phi_0\right)"

In our case:

a. Amplitude

"A=2.0\rm \: m"

b. Frequency

"f=\omega/(2\pi)=0.64\:\rm Hz"

c. Wavelength

"\lambda=2\pi/k=2.1\:\rm m"

d. Speed

"v=f\lambda=0.64*2.1=1.3\:\rm m/s"

e. Initial height at x=1.0 m

"y(1.0, 0) = 2.0 \sin(3.0*1.0 – 4.0*0 + π/2)\\=-1.3\:\rm m"